The capital cost (CC), now, when a family want to have $14,904, per year, starting in year 19 and continuing forever, is close to. Assume rate of return14%
Answer #1

Solution: -

Capital cost =

Value at year 18 = CF/rate = 14904/0.14 = $106457.143

Correct value of it = Value at year 18/(1+r) 18 =

= $106457.143/(1+0.14) 18

= $106457.143 * 0.05083 = $5411.26